# Text-to-Vision Generation with Scene Graph Programming

### v2

@@ Abstract
Generative models like DALL-E and Sora have gained attention by producing implausible images, such as “astronauts riding a horse in space.” Despite the proliferation of text-to-vision models that have inundated the internet with synthetic visuals, from images to 3D assets, current benchmarks predominantly evaluate these models on real-world scenes paired with captions. We introduce Generate Any Scene, a framework that systematically enumerates scene graphs representing a vast array of visual scenes, spanning realistic to imaginative compositions. Generate Any Scene leverages ‘scene graph programming,’ a method for dynamically constructing scene graphs of varying complexity from a structured taxonomy of visual elements. This taxonomy includes numerous objects, attributes, and relations, enabling the synthesis of an almost infinite variety of scene graphs. Using these structured representations, Generate Any Scene translates each scene graph into a caption, enabling scalable evaluation of text-to-vision models through standard metrics. We conduct extensive evaluations across multiple text-to-image, text-to-video, and text-to-3D models, presenting key findings on model performance. We find that DiT-backbone text-to-image models align more closely with input captions than UNet-backbone models. Text-to-video models struggle with balancing dynamics and consistency, while both text-to-video and text-to-3D models show notable gaps in human preference alignment. Additionally, we demonstrate the effectiveness of Generate Any Scene by conducting three practical applications leveraging captions generated by Generate Any Scene: (1) a self-improving framework where models iteratively enhance their performance using generated data, (2) a distillation process to transfer specific strengths from proprietary models to open-source counterparts, and (3) improvements in content moderation by identifying and generating challenging synthetic data.
@@ 1 Introduction
Figure 1: Overview of applications with Generate Any Scene captions: Application 1: (Self-improving): Iteratively enhances a model by generating images with Generate Any Scene captions, selecting the best, and fine-tuning, yielding a performance boost. Application 2: (Distilling limitations): Distills strengths from proprietary models, such as better compositionality and hard concept understanding, into open-source models. Application 3: (Generated content detector): Robustify AI-generated content detection by training on diverse synthetic data generated by Generate Any Scene’s captions.
Artist Marc Chagall said “Great art picks up where nature ends.” The charm of visual content generation lies in the realm of imagination. Since their launch, Dall-E [55, 5] and Sora [7] have promoted their products with implausible generated images of “astronauts riding a horse in space” and “cats playing chess”. With the proliferation of text-to-vision generation models, the internet is now flooded with generated visual content—images, videos, and 3D assets—most generated from user-provided captions [55, 5, 7]. While there are numerous benchmarks designed for evaluating these text-to-vision models, they are typically collections of real-world visual content paired with captions [9, 34, 70]. To quote Marc Chagall again, “If I create from heart, nearly everything works; if from the head, almost nothing.” There is a need for evaluation benchmarks that go beyond real-world scenes and evaluate how well generative models can represent the entire space of imaginary scenes.
@@
Such a comprehensive evaluation requires that we first define the space of the visual content. A long list of prior work [32, 27, 47, 28, 26] has argued that scene graphs [32] are a cognitively grounded [6] representation of the visual space. A scene graph represents objects in a scene as individual nodes in a graph. Each object is modified by attributes, which describe its properties. For example, attributes can describe the material, color, size, and location of the object in the scene. Finally, relationships are edges that connect the nodes. They define the spatial, functional, social, and interactions between objects [42]. For example, in a living room scene, a “table” node might have attributes like “wooden” or “rectangular” and be connected to a “lamp” node through a relation: “on top of.” This systematic scene graph structure provides simple yet effective ways to define and model the scene. Make it an ideal structure for Generate Any Scene to systematically define the diverse space of the visual scenes.
We introduce Generate Any Scene, a system capable of efficiently enumerating the space of scene graphs representing a wide range of visual scenes, from realistic to highly imaginative. Generate Any Scene is powered by what we call scene graph programming, a programmatic approach for composing scene graphs of any complexity using a rich taxonomy of visual elements, and for translating each scene graph into a caption. With a space of synthetically diverse captions, we use Generate Any Scene to prompt Text-to-Vision generation models and evaluate their generations. Like any other representation, scene graphs are also limited: they don’t represent tertiary relationships (e.g. “three people playing frisbee”). Nonetheless, they account for a large space of possibilities. To systematically define and scalably explore the space of user captions, we adopt the scene graph representation [32] to comprehensively evaluate and improve text-to-vision models.
@@
We construct a rich taxonomy of visual concepts consisting of $28,787$ objects, $1,494$ attributes, $10,492$ relations, $2,193$ image/video/3D scene attributes from various sources. Based on these assets, Generate Any Scene can programmatically synthesize an almost infinite number of scene graphs of varying complexity [81]. Besides, Generate Any Scene allows configurable scene graph generation. For example, evaluators can specify the complexity level of the scene graph to be generated or provide a seed scene graph to be expanded. Given an initial scene graph, Generate Any Scene programmatically translates it into a caption, which, when combined with existing text-to-vision metrics, e.g., Clip Score [54] and VQA Score [39], can be used to evaluate any text-to-vision model [62]. By automating these steps, our system ensures both scalability and adaptability, providing researchers and developers with diverse, richly detailed scene graphs and corresponding captions tailored to their specific needs.
With Generate Any Scene’s programmatic generation capability, we release a dataset featuring $10$ million diverse and compositional captions, each paired with a corresponding scene graph. This extensive dataset spans a wide range of visual scenarios, from realistic to highly imaginative compositions, providing an invaluable resource for researchers and practitioners in the Text-to-Vision generation field. We also conduct extensive evaluations of $12$ text-to-image, $9$ text-to-video and $5$ text-to-3D models across a broad spectrum of visual scenes. We have several crucial findings: (1) DiT-backbone models show superior faithfulness and comprehensiveness to input captions than UNet-backbone models, with human-alignment data training helping to bridge some of these gaps. (2) Text-to-Video generation face challenges in balancing dynamics and consistency. (3) All Text-to-Video and Text-to-3D models we evaluate show negative ImageReward Score scores, highlighting a substantial gap in human preference alignment.
Further, we demonstrate the effectiveness of Generate Any Scene by conducting three practical applications leveraging captions generated by Generate Any Scene (Figure 1):
@@
Application 1: Self-improving. We show that our diverse captions can facilitate a framework to iteratively improve Text-to-Vision generation models using their own generations. Given a model, we generate multiple images, identify the highest-scoring one, and use it as new fine-tuning data to improve the model itself. We fine-tune Stable Diffusion v1-5 and achieve an average of 5% performance boost compared with original models, and this method is even better than fine-tuning with the same amount of real images and captions from the Conceptual Captions CC3M [9] over different benchmarks.
@@
Application 2: Distilling limitations. Using our evaluations, we identify limitations in open-sourced models that their proprietary counterparts excel at. Next, we distill these specific capabilities from proprietary models. For example, DaLL-E 3 excels particularly in generating composite images with multiple parts. We distill this capability into Stable Diffusion v1-5, effectively bridging the gap between DaLL-E 3 and Stable Diffusion v1-5.
Application 3: Generated content detector. Content moderation is a vital application, especially as Text-to-Vision generation models improve. We identify which kinds of data content moderation models are bad at detecting, generate more of such content, and retrain the detectors. We train a ViT-T with our generated data and boost its detection capabilities across benchmarks.
2 Related work
Text-to-Vision generation models.
Recent Text-to-Image generation advances are driven by diffusion models that enhance visual fidelity and semantic alignment. Some open-source models [56, 52, 35, 51, 1] use UNet backbones to refine images iteratively. In parallel, Diffusion Transformers (DiTs) architectures[16, 11, 12, 33] have emerged as a better alternative in capturing long-range dependencies and improving coherence. Proprietary models like DALL-E 3 [5] and Imagen 3 [2] still set the state-of-the-art. Based on Text-to-Image generation method, Text-to-Video generation models typically utilize time-aware architectures to ensure temporal coherence across frames [19, 65, 30, 67, 74, 10, 80, 84]. In Text-to-3D generation, recent proposed models [69, 53, 66, 44, 38] integrate the diffusion models with Neural Radiance Fields (NeRF) rendering to generate diverse 3D object. In this work, we systematically evaluate and deeply analyze these Text-to-Vision generation models.
Synthetic prompts for Text-to-Vision generation.
Prompts for Text-to-Vision generation models vary greatly in diversity, complexity, and compositionality. This variation makes it challenging and costly to collect large-scale and diverse prompts written by humans. Consequently, synthetic prompts have been widely used for both training [37, 62, 36, 83, 43, 63, 48, 71, 76, 82] and evaluation purposes [23]. For example, training methods like LLM-Grounded Diffusion [37] leverage LLM-generated synthetic text prompts to enhance the model’s understanding and alignment with human instruction. For evaluation, benchmarks such as T2I-CompBench [23] and T2V-CompBench [63] utilize benchmarks generated by LLMs. However, while LLMs can generate large-scale natural prompts, controlling their content is difficult, and they might exhibit systematic bias. In this work, we propose a programmatic scene graph-based prompt generation system that can generate infinitely diverse scene graph-based prompts for evaluating and improving Text-to-Vision generation models.
Finetuning techniques for Text-to-Vision generation.
To accommodate the diverse applications and personalization needs in text-to-vision models, numerous fine-tuning techniques have been developed. For example, LoRA [21] reduces the computational resources required for fine-tuning by approximating weight updates with low-rank matrices, enabling efficient adaptation to new tasks. Textual Inversion [46, 17] introduces new word embeddings to represent novel concepts, allowing models to generate images of user-specified content without extensively altering the original parameters. DreamBooth [57] fine-tunes models on a small set of personalized images to capture specific subjects or styles, facilitating customized content generation. DreamSync [62], provides a novel way to enable models to learn from the high-quality output generated by itself. In this work, we use several fine-tuning techniques with our programmatic scene graph-based prompts to improve Text-to-Vision generation models.
@@ 3 Generate Any Scene
Figure 2: The generation pipeline of Generate Any Scene: Step 1: The system enumerates scene graph structures that contain objects, attributes, and relations based on complexity, and queries the corresponding scene graph structure that satisfies the needs. Step 2: It populates these structures with metadata, assigning specific content to each node. Scene graphs are completed in this step. Step 3: In addition to the scene graph, scene attributes—such as art style and camera settings—are sampled to provide contextual depth beyond the scene graph. Step 4: The Generate Any Scene system combines the scene graph and scene attributes, such as art style and camera settings, and then translates them into a coherent caption by organizing the elements into structured text.
We present our implementation of Generate Any Scene system. (Figure 2) It programmatically synthesizes diverse scene graphs in terms of both structure and content and translates them into corresponding captions.
Scene graph. A scene graph is a structured representation of a visual scene, where objects are represented as nodes, their attributes (such as color and shape) are properties of those nodes, and their relationships (such as spatial or semantic connections) are represented as edges. In recent years, scene graphs have played a crucial role in visual understanding tasks, such as those found in Visual Genome [32] and GQA [25] for visual question answering (VQA). Their utility has expanded to various Text-to-Vision generation tasks. For example, the DSG score [13] leverages MLMs to evaluate how well captions align with generated scenes by analyzing scene graphs.
@@
Objects 28,787 WordNet [45]
Attributes 1,494 Wikipedia [72], etc.
Relations 10,492 Robin [49]
Scene Attributes 2,193 Places365 [41], etc.
Table 1: Summary of the quantities and source of visual elements.
Taxonomy of visual elements. To construct a scene graph, we use three main metadata types: objects, attributes, and relations. We also have scene attributes that capture the board aspect of the caption, such as art style, to create a complete visual caption. The numbers and the source of our metadata are illustrated in Table 1. Additionally, we build a taxonomy that categorizes metadata into distinct levels and types, enabling fine-grained analysis. This structure allows for detailed assessments, such as evaluating model performance on “flower” as a general concept and on specific subcategories like “daisy.” More details in Appendix C.
3.1 Scene graph programming
Step 1: Scene graph structure enumeration and query. Our system first generates and stores a variety of scene graph structures based on a specified level of complexity, defined by the total number of objects, relationships, and attributes in each graph. The process begins by determining the number of object nodes, and then by systematically enumerating different combinations of relationships among these objects and their associated attributes. Once all graph structures meeting the complexity constraint are enumerated, they are stored in a database for later use. This enumeration process is executed only once for each level of complexity, allowing us to efficiently query the database for suitable templates when needed.
@@ 3.1 Scene graph programming
Step 2: Populate the scene graph structure with metadata. Given a scene graph structure, the next step involves populating the graph with metadata. For each object node, attribute node, and relation edge, we sample the corresponding content from our metadata. This process is highly customizable: users can define the topics and types of metadata to be included (e.g., selecting only common metadata or specifying particular relationships between particular objects, among other options). By determining the scope of metadata sampling, we can precisely control the final content of the captions and easily extend the diversity and richness in the scene graphs by incorporating new datasets.
@@
Step 3: Sampling scene attributes. In addition to scene graphs that capture the visual content of the image, we also include scene attributes that describe aspects such as the art style, viewpoint, time span (for video), and 3D attributes (for 3D content). These scene attributes are sampled directly from our metadata, creating a list that provides contextual details to enrich the description of the visual content.
Step 4: Translate scene graph to caption. We introduce an algorithm that converts scene graphs and a list of scene attributes into captions. The algorithm processes the scene graph in topological order, transforming each object, its attributes, and relational edges into descriptive text. To maintain coherence, it tracks each concept’s occurrence, distinguishing objects with identical names using terms like “the first” or “the second.” Objects that have been previously referenced without new relations are skipped to avoid misreferencing. This approach enhances caption clarity by preventing repetition and maintaining a logical reference.
4 Evaluating Text-to-Vision generation models
4.1 Experiment Settings
Details of experiment settings are in Appendix D.
Models. We conduct experiments on 12 Text-to-image models [1, 52, 56, 35, 51, 16, 11, 12, 33, 5], 9 Text-to-Video models [67, 61, 30, 19, 65, 74, 84, 80, 10], and 5 Text-to-3D models [53, 38, 66, 69, 44]. Text-to-image models are evaluated at a resolution of 1024 × 1024 pixels. We standardize the frame length to 16 across all Text-to-Video models for fair comparisons. For Text-to-3D, we generate videos by rendering from 120 viewpoints.
Metrics. Across all Text-to-Vision generation tasks, we use Clip Score [8] (semantic similarity), VQA Score [39] (faithfulness), TIFA Score [22, 13] (faithfulness), Pick Score [31] (human preference), and ImageReward Score [77] (human preference) as general metrics, and for Text-to-Video generation, VBench [24] for fine-grained video analysis like consistency and dynamics.
Synthetic captions. We evaluate our Text-to-Image generation and Text-to-Video generation models on 10K randomly generated captions, with scene graph complexity ranging from 3 to 12 and scene attributes from 0 to 5, using unrestricted metadata. For Text-to-3D generation models, due to their limitations in handling complex captions and time-intensive generation, we restrict scene graph complexity to 1-3, scene attributes to 0-2, and evaluate on 1K captions.
4.2 Overall results
We evaluate Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation models on Generate Any Scene. Here, we only list key findings; more details and raw results can be found in Appendix D.
Text-to-Image generation results. (Figure 3)
  1. 1.
  2. 2.
  3. 3.
    The closed-source model DaLL-E 3 maintains a significant lead in VQA Score, TIFA Score, and ImageReward Score, demonstrating strong faithfulness and alignment with prompts across generated content.
Text-to-Video generation results. (Table 2,3)
Model clip score pick score image reward score VQA score TiFA score
VideoCraft2 [10] 0.2398 0.1976 -0.4202 0.5018 0.2466
AnimateLCM [65] 0.2450 0.1987 -0.5754 0.4816 0.2176
AnimateDiff [19] 0.2610 0.1959 -0.7301 0.5255 0.2208
Open-Sora 1.2 [84] 0.2259 0.1928 -0.6277 0.5519 0.2414
FreeInit [74] 0.2579 0.1950 -0.9335 0.5123 0.2047
ModelScope [67] 0.2041 0.1886 -1.9172 0.3840 0.1219
Text2Video-Zero [30] 0.2539 0.1933 -1.2050 0.4753 0.1952
CogVideoX-2B [80] 0.2038 0.1901 -1.2301 0.4585 0.1997
ZeroScope [61] 0.2289 0.1933 -1.1599 0.4892 0.2388
Table 2: Overall performance of Text-to-Video generation models over 10K Generate Any Scene captions.Red Cell is the highest score. Yellow Cell is the second highest score.
Model subject consistency background consistency motion smoothness dynamic degree
Open-Sora 1.2 0.9964 0.9907 0.9973 0.0044
Text2Video-Zero 0.8471 0.9030 0.8301 0.9999
VideoCraft2 0.9768 0.9688 0.9833 0.3556
AnimateDiff 0.9823 0.9733 0.9859 0.1406
FreeInit 0.9581 0.9571 0.9752 0.4440
ModelScope 0.9795 0.9831 0.9803 0.1281
AnimateLCM 0.9883 0.9802 0.9887 0.0612
CogVideoX-2B 0.9583 0.9602 0.9823 0.4980
ZeroScope 0.9814 0.9811 0.9919 0.1670
Table 3: Overall performance of Text-to-Video generation models over 10K Generate Any Scene captions with VBench metrics. Red Cell is the highest score. Blue Cell is the lowest score.
  1. 1.
    Text-to-video models face challenges in balancing dynamics and consistency (Table 3). This is especially evident in Open-Sora 1.2, which achieves high consistency but minimal dynamics, and Text2Video-Zero, which excels in dynamics but suffers from frame inconsistency.
  2. 2.
  3. 3.
    VideoCrafter2 strikes a balance across key metrics, leading in human-preference alignment, faithfulness, consistency, and dynamic.
Text-to-3D generation results. (Table 4)
Model clip score pick score vqa score tifa score image reward score
Latent-NeRF [44] 0.2115 0.1910 0.4767 0.2216 -1.5311
DreamFusion-sd [53] 0.1961 0.1906 0.4421 0.1657 -1.5582
Magic3D-sd [38] 0.1947 0.1903 0.4193 0.1537 -1.6327
SJC [66] 0.2191 0.1915 0.5015 0.2563 -1.4370
DreamFusion-IF [53] 0.1828 0.1857 0.3872 0.1416 -1.9353
Magic3D-IF [38] 0.1919 0.1866 0.4039 0.1537 -1.8465
ProlificDreamer [69] 0.2125 0.1940 0.5411 0.2704 -1.2774
Table 4: Overall performance of Text-to-3D generation models over 10K Generate Any Scene captions.
  1. 1.
    ProlificDreamer outperforms other models, particularly in ImageReward Score, VQA Score and TIFA Score.
  2. 2.
5 Application 1: Self-Improving Models

### v3

# Ziqi Gao [^1] footnotemark: 1, Weikai Huang [^1] footnotemark: 1, Jieyu Zhang1, Aniruddha Kembhavi2, Ranjay Krishna1,2

@@ Abstract
Recent advances in text-to-vision generation excel in visual fidelity but struggle with compositional generalization and semantic alignment. Existing datasets are noisy and weakly compositional, limiting models’ understanding of complex scenes, while scalable solutions for dense, high-quality annotations remain a challenge. We introduce Generate Any Scene, a data engine that systematically enumerates scene graphs representing the combinatorial array of possible visual scenes. Generate Any Scene dynamically constructs scene graphs of varying complexity from a structured taxonomy of objects, attributes, and relations. Given a sampled scene graph, Generate Any Scene translates it into a caption for text-to-image or text-to-video generation; it also translates it into a set of visual question answers that allow automatic evaluation and reward modeling of semantic alignment. Using Generate Any Scene, we first design a self-improving framework where models iteratively enhance their performance using generated data. SDv1.5 achieves an average 4% improvement over baselines and surpassing fine-tuning on CC3M. Second, we also design a distillation algorithm to transfer specific strengths from proprietary models to their open-source counterparts. Using fewer than 800 synthetic captions, we fine-tune SDv1.5 and achieve a 10% increase in TIFA score on compositional and hard concept generation. Third, we create a reward model to align model generation with semantic accuracy at a low cost. Using GRPO algorithm, we fine-tune SimpleAR-0.5B-SFT and surpass CLIP-based methods by +5% on DPG-Bench. Finally, we apply these ideas to the downstream task of content moderation where we train models to identify challenging cases by learning from synthetic data.
@@ 1 Introduction
Despite the high-fidelity of modern generative models (text-to-image and text-to-video), we are yet to witness wide-spread adoption (Brooks et al., 2024; Ramesh et al., 2022; Betker et al., 2023; Wang et al., 2024c; Chen et al., 2023b). Controllability remains out of reach (Peng et al., 2024). Generated content appears realistic but often falls short of semantic alignment (Huang et al., 2023; Sun et al., 2024; Ghosh et al., 2023; Hu et al., 2024a). Users prompt models with a specific concept in mind. For example, when prompted to generate a scene of a “A black dog chasing after a rabbit that is eating the grass, in Van Gogh’s style, with starlight lightening”, some models are likely to generate an image of a dog but might miss the rabbit or get the style incorrect.
We hypothesize that these limitations stem not only from architectural bottlenecks but more fundamentally from the lack of structured, compositionally rich training data (Betker et al., 2023), especially those with uncommon compositions. Popular datasets such as LAION (Schuhmann et al., 2022) and CC3M (Sharma et al., 2018) predominantly consist of web-crawled image-caption pairs, which are inherently noisy, weakly compositional, and biased toward single-object, coarse-grained descriptions. Such datasets lack explicit grounding of object-attribute relations and multi-object interactions, restricting models’ ability to generalize to complex visual scenes. Efforts to enhance caption quality (Betker et al., 2023; Li et al., 2024d) have demonstrated that enhancing the compositional density and semantic richness of captions can significantly improve generative performance. Nevertheless, manual curation of such dense compositional annotations is labor-intensive, while automatic annotation methods (e.g., via MLMs) suffer from hallucination and semantic noise.
@@
Constructing a compositional dataset requires that we first define the space of the visual content. Scene graphs are one such representation of the visual space (Krishna et al., 2017; Johnson et al., 2015; Ost et al., 2021; Johnson et al., 2018; Ji et al., 2020), grounded in cognitive science (Biederman, 1987). A scene graph represents objects in a scene as individual nodes in a graph. Each object is modified by attributes, which describe its properties. For example, attributes can describe the material, color, size, and location of the object in the scene. Finally, relationships are edges that connect the nodes. They define the spatial, functional, social, and interactions between objects (Lu et al., 2016). For example, in a living room scene, a “table” node might have attributes like “wooden” or “rectangular” and be connected to a “lamp” node through a relation: “on top of”. This systematic scene graph structure provides simple yet effective ways to define and model the scene. As such, scene graphs are an ideal foundation for systematically defining the compositional space of visual content in text-to-vision generation.
@@
We introduce Generate Any Scene, a system capable of efficiently enumerating the space of scene graphs representing a wide range of visual scenes. Generate Any Scene composes scene graphs of any structure using a rich taxonomy of visual elements, translating each scene graph into an input caption and visual question answers to evaluate the output image or video. In particular, we first construct a rich taxonomy of visual concepts consisting of $28,787$ objects, $1,494$ attributes, $10,492$ relations, $2,193$ scene attributes from various sources. Based on these assets, Generate Any Scene can synthesize an almost infinite number of scene graphs of varying complexity (Zhang et al., 2024). Besides, Generate Any Scene allows configurable scene graph generation. For example, evaluators can specify the complexity level of the scene graph to be generated or provide a seed scene graph to be expanded. By automating these steps, our system ensures both scalability and adaptability, providing researchers and developers with diverse, richly detailed scene graphs and corresponding captions tailored to their specific needs. We also conduct comprehensive text-to-vision evaluations using our generated captions, as detailed in Appendix A.
@@
We show that Generate Any Scene can allow generation models to self-improve. Our diverse captions can facilitate a framework to iteratively improve Text-to-Vision generation models using their own generations. Given a model, we generate multiple images, identify the highest-scoring one, and use it as new fine-tuning data to improve the model itself. We fine-tune SDv1.5 (Rombach et al., 2022b) and achieve an average of 4% performance boost compared with original models, and this method is even better than fine-tuning with the same amount of real images and captions from the Conceptual Captions CC3M over different benchmarks.
@@
We also use Generate Any Scene to design targeted distillation algorithms. Using our evaluations, we identify limitations in open-sourced models that their proprietary counterparts excel at. Next, we distill these specific capabilities from proprietary models. For example, DaLL-E 3 (Betker et al., 2023) excels particularly in generating composite images with multiple parts. We distill this capability into SDv1.5, effectively bridging the gap between DaLL-E 3 and SDv1.5. After targeted fine-tuning, SDv1.5 achieves a 10% increase in TIFA score (Hu et al., 2023) for compositional tasks and hard concept generation.
Then we propose a low-cost scene graph-based reward model for RLHF (Ouyang et al., 2022) in text-to-image generation. By leveraging synthetic scene graphs generated by Generate Any Scene, we generate exhaustive question-answer pairs that cover all objects, attributes, and relationships in the caption. Our method enables fine-grained, compositional reward modeling without manual annotation or heavy LLM inference. With GRPO (Shao et al., 2024), we fine-tune SimpleAR-0.5B-SFT (Wang et al., 2025) using a scene graph reward model, achieving better compositional alignment than CLIP-based methods (Radford et al., 2021) (+5% on DPG-Bench (Hu et al., 2024b)).
Finally, we apply Generate Any Scene to the downstream application of content moderation. Content moderation is a vital application, especially as Text-to-Vision generation models improve. A key challenge lies in the limited diversity of existing training data. To address this, we leverage Generate Any Scene to generate diverse and compositional captions, creating synthetic training data that complements existing datasets. By retraining a ViT-T (Dosovitskiy et al., 2021) detector with our enriched dataset, we enhance its detection performance, particularly in cross-model and cross-dataset scenarios.
@@ 2 Generate Any Scene
In this section, we present Generate Any Scene (Figure 1), a data engine that systematically synthesizes diverse scene graphs in terms of both structure and content and translates them into corresponding captions.
Scene graph. A scene graph is a structured representation of a visual scene, where objects are represented as nodes, their attributes (such as color and shape) are properties of those nodes, and the relationships between objects (such as spatial or semantic connections) are represented as edges. In recent years, scene graphs have played a crucial role in visual understanding tasks, such as those found in Visual Genome Krishna et al. (2017) and GQA Hudson & Manning (2019) for visual question answering (VQA). Their utility has expanded to various Text-to-Vision generation tasks. For example, the DSG Cho et al. (2023) and DPG Hu et al. (2024a) benchmarks leverage scene graphs to evaluate how well generated images align with captions.
Taxonomy of visual elements. To construct a scene graph, we use three main metadata types: objects, attributes, and relations. We further introduce scene attributes that capture global visual contexts, such as art style, to facilitate comprehensive caption synthesis. The statistics and source of our metadata are shown in Table 1. Additionally, we build a hierarchical taxonomy that categorizes metadata into distinct levels and types, enabling fine-grained analysis. This structure supports precise content synthesis, from broad concepts like “flower” to fine-grained instances such as “daisy.”
Table 1: Summary of the quantities and sources of visual elements. Details are in Appendix B.
@@
Objects 28,787 WordNet (Miller, 1995)
Attributes 1,494 Wikipedia (Wikipedia Contributors, 2024), etc.
Relations 10,492 Synthetic Visual Genome (Park et al., 2025)
Scene Attributes 2,193 Places365 (López-Cifuentes et al., 2020), etc.
2.1 Generating data with scene graphs
Step 1: Scene graph structure enumeration. Our engine pre-computes a library of directed scene-graph topologies subject to user-specified structural constraints: complexity (total number of objects, relations, and attributes) (Grunde-McLaughlin et al., 2021), average node degree, and number of connected components. We first sample the number of object nodes and then systematically enumerate feasible edge sets and attribute attachments that satisfy these constraints. We provide 3 optional controls: (i) degree-profile bounds per-node in/out-degree, (ii) seed-graph preservation embeds a user-provided seed graph as a subgraph of each enumerated structure, and (3) commonsense plausibility filtering prunes implausible contents while retaining compositional diversity. All enumerations are performed once per parameter tuple and cached for fast querying.
@@ 2.1 Generating data with scene graphs
Step 2: Populate the scene graph structure with metadata. Given a generated scene graph structure, the next step involves populating the graph with metadata. For each object node, attribute node, and relation edge, we sample the corresponding content from our metadata. This process is highly customizable and controllable: users can define the topics and types of metadata to include, for instance, by selecting only commonsense metadata or specifying relationships between particular objects. By determining the scope of metadata sampling, we can precisely control the final content of the captions and easily extend the diversity and richness of scene graphs by adding new metadata.
@@
Step 3: Sample scene attributes. We also include scene attributes that describe aspects such as the art style, viewpoint, time span (for video), and 3D attributes (for 3D content). These scene attributes are sampled directly from our metadata, creating a list that provides contextual details to enrich the description of the visual content.
Step 4: Translate scene graph to caption. We introduce a deterministic and programmatic algorithm that converts scene graphs with scene attributes into captions. It traverses scene graphs by converting objects/attributes/relations into descriptive text in topological order, while tracking each object’s references to ensure coherence. Programmatic grammar rules are employed (e.g., disambiguating identical objects with “the first/second” and skipping already mentioned objects) to prevent duplication and misreference, resulting in clear captions. We also provide LLM paraphrasing as an optional step to diversify wording; however, our studies (see Appendix A.3) show that paraphrasing does not materially affect results. We adopt the programmatic caption converter as the default for its speed and low hallucination rate.
Step 5: Convert scene graph to a series of question-answer pairs. Given a synthetic scene graph, Generate Any Scene automatically enumerates exhaustive QA pairs using templates that query object attributes (e.g., What color is the sphere?), spatial relations (e.g., What is to the left of the cube?), and other compositional elements. Each answer maps directly to an object, attribute, or edge, ensuring full coverage of the graph at minimal cost. This enables both VQA-based evaluation of generated images and the construction of fine-grained reward models without manual labeling or costly LLM inference.
Figure 1: The generation pipeline of Generate Any Scene. Step 1: Enumerate diverse scene graph structures under user-defined constraints. Step 2: Populate structures with sampled objects, attributes, and relations. Step 3: Sample scene attributes such as style, perspective, or time span. Step 4: Translate scene graph and attributes into coherent captions. Step 5: Automatically generate QA pairs covering all elements for evaluation and reward modeling.
3 Self-Improving models with synthetic captions